OpenStudy (anonymous):

I am really lost in Algebra, I need help(please) finding the slope , if possible, of the line passing througgh each pair of points (2,-1) and 96,3)

jimthompson5910 (jim_thompson5910):

did you mean (2,-1) and (6,3)

jimthompson5910 (jim_thompson5910):

if so, then m = (y2-y1)/(x2-x1) m = (3-(-1))/(6-2) m = (3+1)/(6-2) I'll let you finish. Don't forget to reduce fully.

OpenStudy (anonymous):

yes I did Thank you for the help

jimthompson5910 (jim_thompson5910):

what did you get

OpenStudy (anonymous):

im am working on it now. :(

jimthompson5910 (jim_thompson5910):

that's ok

jimthompson5910 (jim_thompson5910):

take your time

OpenStudy (anonymous):

is 1 right?

jimthompson5910 (jim_thompson5910):

yep

OpenStudy (anonymous):

Yay!!!! Thanks

jimthompson5910 (jim_thompson5910):

yw